OpenStudy (anonymous):

What number should be added to the expression x^2 + 2x to change it into a perfect square trinomial?

OpenStudy (ranga):

To make ax^2 + bx into a perfect square you need to add the term {b / (2a) }^2

OpenStudy (anonymous):

thank you what about this question Find the solution to the following equation by transforming it into a perfect square trinomial. x^2 + 2x = 8

OpenStudy (ranga):

To make x^2 + 2x a perfect square we need to add the term { 2 / (2*1) }^2 which equals 1. Add 1 to both sides: x^2 + 2x = 8 add 1 to both sides: x^2 + 2x + 1 = 8 + 1 (x + 1)^2 = 9 Take square root on both sides: x + 1 = +/- 3 x = +3 - 1 or x = -3 - 1 x = 2 or x = -4

OpenStudy (anonymous):

thank you